Solve ny graphing: x+y=2 2x-y=7
Answer by mananth(12270)

x+y=2 ......................1
2x-y=7 .....................2
Generate y co-ordinates for various values of x and plot both lines.
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 2 ,( 0 , 2 )
when y = 0 x= 2 ,( 2 , 0 )
Equation 2
when x= 0 y= 7 ( 0 , -7 )
when y = 0 x= 3.5 ( 3.5,0 )



(3,-1) is the solution